Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
Nenci - 15:12 Thursday 26 March 2020 (48813) Print this report
Operator Report - Morning shift

> Operator from remote <

ITF found in Science Mode, unfortunately it unlocked at 7:00 UTC and I started to have problems to reach again the Science Mode collecting several unlock (at different nodes) during the acquisition. Finally the Science Mode has been achieved again at 8:18 UTC. The Commissionig activities started as planned at 9:34 UTC, moreover, before start the acoustic injections in CEB, Antonio P. e Marco C. performed a check at the NE building in order to find the origin of the excess of noise reported on the entry #48809, they discovered that the air flux in the cleaning room has been remained switched ON since the last maintenance, standard working condition restored (plot n1 attached).

from 9:34 UTC to 12:06 UTC - COMMISSIONING MODE in order to perform the ENV scheduled activities.

14:13 UTC - ITF in Science Mode at the end of the shift.

EnvMoni
ENV_NE_ACC_Z red flag on DMS. Excess of noise. Problem solved, air flux in the cleaning room switched OFF.

Images attached to this report
Search Help
×

Warning

×